OpenStudy (anonymous):

Evaluate the expression below when x = 5. 3x(2x - 7)

OpenStudy (anonymous):

Plug in x = 5 3x(2x - 7) = 3*5(2*5 -7 ) = 15(10-7) = 15*3=45

OpenStudy (anonymous):

6x^2 -7x =6(25)-7(5) =115